OpenStudy (konradzuse):

Using the standard Normal distribution tables, the area under the standard Normal curve corresponding to Z > –1.62 is A. 0.0044. B. 0.0526. C. 0.9474. D. 0.9956.

OpenStudy (konradzuse):

I get 0.9474

OpenStudy (konradzuse):

1 - 0.9474 = 0.0526 so B?

OpenStudy (konradzuse):

@lgbasallote

jimthompson5910 (jim_thompson5910):

looks good to me, you got it

OpenStudy (konradzuse):

:)

jimthompson5910 (jim_thompson5910):

oh wait, the table represents positive values of z and not negative values (just noticing this) So the answer is actually 0.9474

OpenStudy (konradzuse):

look @ bottom it shows how to get negs?

OpenStudy (konradzuse):

seems like to get a neg yo do 1- #

jimthompson5910 (jim_thompson5910):

This table computes the area to the left of z = k (where k is positive) Since the standard normal distribution is symmetric about 0, this means P(Z > -1.62) = P(Z < 1.62) which shows us that the answer is 0.9474

OpenStudy (konradzuse):

1.62 = 9474 check the bottom it explains it.

jimthompson5910 (jim_thompson5910):

A useful trick to remember is that P(Z > -k) = P(Z < k) for some number k

OpenStudy (konradzuse):

z 9750 = 1.96 z 250 = -1.96 :p

jimthompson5910 (jim_thompson5910):

I see that, but the answer is still 0.9474

OpenStudy (konradzuse):

>(?

jimthompson5910 (jim_thompson5910):

wolfram even confirms it http://www.wolframalpha.com/input/?i=normalcdf%28-1.62%2C10%29

OpenStudy (konradzuse):

look 1 above it... 0.526

OpenStudy (konradzuse):

since it's Z> then I gess yo'r right...

OpenStudy (konradzuse):

>(

jimthompson5910 (jim_thompson5910):

no, the answer you want is in the last row of that table in wolfram alpha

OpenStudy (konradzuse):

so if it said z = 0.250 wold I have been correct?

jimthompson5910 (jim_thompson5910):

where it says -1.62 < z < 10

jimthompson5910 (jim_thompson5910):

you mean P( Z < 0.250) ?

OpenStudy (konradzuse):

I see this is tricky...

OpenStudy (konradzuse):

I'm konfused what the table is showing 4 the neg? z 9750 = 1.96 z 250 = -1.96 :??? When do we use this?

OpenStudy (konradzuse):

Using the standard normal distribution tables, the area under the standard Normal curve corresponding to –0.5 < Z < 1.2 is A. 0.3085. B. 0.8849. C. 0.5764. D. 0.2815. this is actally the next q :)

jimthompson5910 (jim_thompson5910):

the 0.9750 is the area to the left of the curve, the 1.96 is the z-score so z 9750 = 1.96 means P(Z < 1.96) = 0.9750)

OpenStudy (konradzuse):

and what abot the neg? part? -1.96

jimthompson5910 (jim_thompson5910):

if you have a negative, you can flip it so to speak, by using the idea that this distribution is symmetrical about 0 So that's why they're subtracting 0.9750 from 1 to get 0.0250

jimthompson5910 (jim_thompson5910):

The rule they are using is this P(Z > -k) = P(Z < k) for some positive number k

OpenStudy (konradzuse):

yes thats what I was saying....

OpenStudy (konradzuse):

so in this case since they as x<z<y we have to select the pos value?

OpenStudy (konradzuse):

if it wanted z = -1.69 then we cold have bused the other one?

OpenStudy (konradzuse):

or shold I say Z>

jimthompson5910 (jim_thompson5910):

not sure where you're getting -1.69

OpenStudy (konradzuse):

z< -1.96

jimthompson5910 (jim_thompson5910):

oh ok

OpenStudy (konradzuse):

err -1.62

OpenStudy (konradzuse):

they buse 1.96 :P

jimthompson5910 (jim_thompson5910):

if you wanted to find P(Z > -1.62) then you would use the trick above to get P(Z < 1.62), from there, you'd use the table

OpenStudy (konradzuse):

Ah I see how this i snow I had to look at it again.. So Z < 1.62 only ntil -1.62 ho...

OpenStudy (konradzuse):

P(Z < 1.62), wold give the same answer?

jimthompson5910 (jim_thompson5910):

yes, P(Z > -1.62) and P(Z < 1.62) are equivalent

OpenStudy (konradzuse):

I mean on the table 1.62 is 1.62 doesn't matter if it's P(Z < 1.62), or P(Z < -1.62),

jimthompson5910 (jim_thompson5910):

oh, it's positive

OpenStudy (konradzuse):

or does one yeild one half and the other the other half? Like the .975 and .025

jimthompson5910 (jim_thompson5910):

so it's just P(Z < 1.62)

OpenStudy (konradzuse):

it's always one or the other? no Z = 1.62?

OpenStudy (konradzuse):

well if Z > 1.62 we have an isse

jimthompson5910 (jim_thompson5910):

no, you can't compute the area from 1.62 to 1.62...that doesn't make any sense

OpenStudy (konradzuse):

what?

jimthompson5910 (jim_thompson5910):

if you want Z > 1.62, then find Z < 1.62 and subtract that from 1

OpenStudy (konradzuse):

ic. os if it's above or below IC G|< Z < |G it will be 1- #?

OpenStudy (konradzuse):

now we are getting somewhere

OpenStudy (konradzuse):

in this next Q Using the standard normal distribution tables, the area under the standard Normal curve corresponding to –0.5 < Z < 1.2 is A. 0.3085. B. 0.8849. C. 0.5764. D. 0.2815. We need to do it 2x right? once 4 -0.5 and once 4 1.2? :)

OpenStudy (konradzuse):

or Idk since it's 1 answer.

jimthompson5910 (jim_thompson5910):

–0.5 < Z < 1.2 is the same as –0.5 < Z and Z < 1.2

OpenStudy (konradzuse):

I thoght abot sing both on the table.

jimthompson5910 (jim_thompson5910):

which turns into Z > -0.5 and Z < 1.2

OpenStudy (konradzuse):

0.5199 &0.8849

OpenStudy (konradzuse):

0.8849 = B

OpenStudy (konradzuse):

There might be more to this Q tho....

OpenStudy (konradzuse):

actally above 0.5 is

OpenStudy (konradzuse):

0.4801

jimthompson5910 (jim_thompson5910):

Z < 0.5 is 0.6915

jimthompson5910 (jim_thompson5910):

So P( Z > -0.5) = 0.6915

OpenStudy (konradzuse):

oh crap I looked @ 0.05 HAAHAH

OpenStudy (konradzuse):

0.6915 isn't there so is it B?

jimthompson5910 (jim_thompson5910):

Now use the rule P(a < Z < b) = P(Z < b) - P(Z < a)

OpenStudy (konradzuse):

I was thinking that bt it woldonly be .1....?

OpenStudy (konradzuse):

0.1934

jimthompson5910 (jim_thompson5910):

P(a < Z < b) = P(Z < b) - P(Z < a) P(-0.5 < Z < 1.2) = P(Z < 1.2) - P(Z < -0.5) P(-0.5 < Z < 1.2) = P(Z < 1.2) - P(Z > 0.5) P(-0.5 < Z < 1.2) = P(Z < 1.2) - [ 1 - P(Z < 0.5) ] P(-0.5 < Z < 1.2) = 0.8849 - [ 1 - 0.6915 ] P(-0.5 < Z < 1.2) = 0.8849 - [ 0.3085 ] P(-0.5 < Z < 1.2) = 0.5764

OpenStudy (konradzuse):

where does that 1 come bfrom?

OpenStudy (konradzuse):

oic why is itdiff bfrom the prev answer...?

OpenStudy (konradzuse):

oh I see format...

OpenStudy (konradzuse):

thanks :)

OpenStudy (konradzuse):

http://openstudy.com/study#/updates/505cf41be4b0583d5cd15dea next q's!!!! :)

jimthompson5910 (jim_thompson5910):

yeah it's a bit messy, but you at least see all the nitty gritty details

jimthompson5910 (jim_thompson5910):

sry i have to get going

OpenStudy (konradzuse):

yeah I'm glad I learned this.

OpenStudy (konradzuse):

Oh okay >(. I think I'm correct, no one's verifying me :(

jimthompson5910 (jim_thompson5910):

you at least have wolfram alpha to check your work

jimthompson5910 (jim_thompson5910):

do you know how to use it?

OpenStudy (konradzuse):

thanks again!! yeah I'm not sre how

OpenStudy (konradzuse):

calc 2 I doo, not stats..

jimthompson5910 (jim_thompson5910):

type normalcdf(a,b) and replace 'a' and 'b' with any numbers you want (where a < b) Examples: normalcdf(1,2) computes P( 1 < Z < 2) normalcdf(-0.5,1.2) computes P( -0.5 < Z < 1.2) normalcdf(2,10) computes P( 2 < Z < 10) ...which is basically the same as P( Z > 2)

OpenStudy (konradzuse):

what abot mean and standard deviation?

jimthompson5910 (jim_thompson5910):

you can either a) convert each raw score to a standard z score or b) type normalcdf(a,b,mu,sigma) which will compute P( a < X < b) with mean mu and standard deviation sigma if you leave off mu and sigma, wolfram alpha will use the default mu = 0 and sigma = 1

OpenStudy (konradzuse):

okay :) thanks :)

jimthompson5910 (jim_thompson5910):

np